OpenStudy (anonymous):

What is the value of the missing exponent? (3x3)^?=27x^9

OpenStudy (anonymous):

Use the properties of exponents to simplify. \[(3^2)^x = (3^3)^9\]\[3^{2x} = 3^{27}\]\[2x = 27\]\[x = \frac{27}{2}\]